A car of mass 1000 kg is moving at speed of 20 m/s on horizontal surface if the car moves and come to rest after traveling 50m distance then what force of friction tends to stop it ?


1
Expert's answer
2021-05-05T19:05:40-0400

We can find the force of friction from the work-kinetic energy theorem:


"\Delta KE=W,""KE_f-KE_i=-F_{fr}d,""-\dfrac{1}{2}mv_i^2=-F_{fr}d,""F_{fr}=\dfrac{mv_i^2}{2d},""F_{fr}=\dfrac{1000\ kg\cdot(20\ \dfrac{m}{s})^2}{2\cdot50\ m}=4000\ N."